OpenStudy (anonymous):

find the equation of the line that contains the point (8,1) and has slope -5....thanks for the help!

OpenStudy (amistre64):

y = -5(x-8)+1

OpenStudy (amistre64):

y = -5x +41

OpenStudy (anonymous):

do i add -5x to both sides to solve for y?

OpenStudy (anonymous):

its already solved for y. Do you mean solve for x in terms of y?

OpenStudy (anonymous):

I guess I wanted to know if anything further needed to be done or if this was the complete answer. Sorry...I am horrible at math! :)

OpenStudy (amistre64):

there are many ways to express the equation of a line; this is the most common form tho.

OpenStudy (anonymous):

yes the usual is y=mx+c in wich this is already displayed

OpenStudy (amistre64):

5x+y = 41 y = -5x +41 y-1 = -5(x-8) 5x +y -41 = 0

OpenStudy (anonymous):

m being the gradient and c the intercept, usually the way thats most useful

OpenStudy (anonymous):

gotcha...thank you for the help and for the extra examples. I really appreciate it!

OpenStudy (anonymous):

no problem :D

OpenStudy (anonymous):

one more question....would the equation be the same if I have two points rather than a slope? Examples points (0,4) and (1.-3)

OpenStudy (anonymous):

sorry that should be (1,-3)

OpenStudy (anonymous):

The slope is found using diff in y / diff in x

OpenStudy (anonymous):

\[\frac{4-(-3)}{1}\] = -7 so the equation becomes y=-7x+4

OpenStudy (anonymous):

so you subtracted the two y points from each other and the two x points from each other, than divide that to find the slope? so in the equation y represents the slope?

OpenStudy (anonymous):

sorry missed the -1 in the denominator

OpenStudy (anonymous):

The y represents the slope and the intercept y=Mx+C where M is the slope and C is where it crosses the y axis

OpenStudy (anonymous):

So the equation represent the line

OpenStudy (anonymous):

so in its fullness we get \[y= \frac{(4-(-3))}{(0-1)}x + 4\]

OpenStudy (anonymous):

I see it now! :D

OpenStudy (anonymous):

kewl :)

OpenStudy (anonymous):

thank u for ur patience with me!